OpenStudy (anonymous):

Factor. 25b^2 - 36 a. (5b+ 6)(5b- 6) b. (5b+ 9)(5b- 4) c. (5b- 6)(5b- 6)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

25b^2 = (5b)^2 36 = 6^2 now use difference of squares..